# The Angular Momentum Penrose Inequality

### v2

@@ A MOTS is stable if the principal eigenvalue of the MOTS stability operator
    Spectral correspondence: By [8, Proposition 3.4], the cylindrical decay rate satisfies $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}$ for strictly stable MOTS. This relationship arises from the linearized Jang equation at the MOTS.
@@
    The orbit radius vanishes linearly at the poles: $\rho(x)=O(\mathrm{dist}(x,p_{\pm}))$ as $x\to p_{\pm}$.
@@
Step 3: Regularity at the poles. The mean curvature $H$ of $\Sigma$ is finite and smooth everywhere, including at the poles. This is because $\Sigma$ is a smooth embedded surface (by elliptic regularity for the MOTS equation). The apparent singularity in coordinate expressions for $H$ (involving terms like $1/\rho$) is a coordinate artifact that cancels when computed correctly.
@@
However, this term appears in the second fundamental form component $A_{\phi\phi}$, which when traced with the metric involves an additional factor of $1/f^{2}$ from the inverse metric $g^{\phi\phi}=1/f^{2}$. The full expression for the mean curvature contribution from this term is:
@@
This does diverge in coordinates, but the metric $g_{\phi\phi}=f^{2}\to 0$ at the same rate, so the trace $H=g^{ij}A_{ij}$ requires care.
@@
Step 4: Orbit radius scaling. For a smooth axisymmetric surface $\Sigma\cong S^{2}$ intersecting the axis at poles $p_{\pm}$, the orbit radius $\rho=|\eta|$ vanishes linearly at the poles. To see this, use geodesic normal coordinates $(s,\theta)$ centered at a pole $p$, where $s$ is geodesic distance from $p$ and $\theta$ is the azimuthal angle. Near $p$:
$\rho(s,\theta)=s\sin(\alpha(s))+O(s^{2})$
where $\alpha(s)$ is the angle between the geodesic and the axis. For a smooth surface intersecting the axis transversally (as any axisymmetric $S^{2}$ must), $\sin(\alpha(0))>0$, giving $\rho=O(s)=O(\mathrm{dist}(x,p))$.
Alternative verification via embedding: For the Kerr horizon, which is topologically $S^{2}$, explicit computation in Boyer-Lindquist coordinates shows $\rho=\sqrt{r_{+}^{2}+a^{2}}\sin\theta$ where $\theta$ is the polar angle. Near the poles ($\theta=0,\pi$), $\rho\sim|\theta|\sim\mathrm{dist}(x,p)$, confirming linear vanishing. ∎
Remark 4.8 (Axis Intersection for MOTS).
For a stable MOTS $\Sigma\cong S^{2}$ in axisymmetric initial data, $\Sigma$ must intersect the symmetry axis $\Gamma$ at exactly two poles. This follows from topological considerations: the Killing field $\eta$ vanishes precisely on $\Gamma$, and any embedded $S^{2}$ symmetric under rotation must contain the fixed points of the $U(1)$ action. The key technical consequence is that the twist perturbation estimates must handle the degenerate case $\rho\to 0$ at the poles—see Lemma 4.9 below.
@@
where $\mathcal{T}_{0}$ involves the twist 1-form $\omega$ contracted with the graph normal. The crucial observation is that $\mathcal{T}$ is proportional to $\rho^{2}$, not merely $\rho$.
Step 2: Axis regularity of the twist. By the axis regularity condition for axisymmetric spacetimes [53, Chapter 7], the twist 1-form $\omega$ satisfies:
@@
Explicit axis regularity conditions for the twist potential $\omega$: The twist 1-form $\omega$ arises from the frame-dragging components of $K$ via the formula $K_{\phi i}=\frac{1}{2}\rho^{2}\omega_{i}$ for $i\in\{r,z\}$ in Weyl–Papapetrou coordinates. The elementary flatness condition at the axis [53, Section 7.1] requires that the spacetime be locally flat on the axis, which imposes:
@@
    Hölder regularity in weighted spaces: The twist 1-form satisfies $\omega\in C^{0,\alpha_{H}}_{\rho}(\mathcal{Q})$, the weighted Hölder space with weight $\rho$. Explicitly:
    $\|\omega\|_{C^{0,\alpha_{H}}_{\rho}}:=\sup_{\mathcal{Q}}|\omega|+\sup_{x\neq y}\frac{|\omega(x)-\omega(y)|}{d(x,y)^{\alpha_{H}}}<\infty.$
@@
    Weighted norms and the axis: The weighted Hölder norm $\|\cdot\|_{C^{k,\alpha_{H}}_{-\tau}}$ (Definition 4.2) involves the radial weight $\langle r\rangle^{-\tau}$ for asymptotic decay, but near the axis we use the $\rho$-weighted regularity $C^{k,\alpha_{H}}_{\rho}$ as in condition (AR3). This hybrid weighting—polynomial in $r$ for asymptotics, $\rho$-scaled for the axis—is standard in the analysis of axisymmetric elliptic problems [19, 23].
@@ 4.3 Axisymmetric Setting
    Existence and uniqueness: The axisymmetric Jang equation admits a solution $f:M\setminus\Sigma\to\mathbb{R}$, unique up to an additive constant. The solution satisfies $f\in C^{2,\alpha}_{\mathrm{loc}}(M\setminus\Sigma)\cap C^{0,1}(M)$ (locally $C^{2,\alpha}$ away from $\Sigma$, globally Lipschitz).
@@
    $f(x)=C_{0}\ln(1/s)+A(y)+R(s,y),\quad C_{0}=\frac{|\theta^{-}|}{2}>0,$
@@
    • $A\in C^{2,\alpha}(\Sigma)$ is a smooth function on $\Sigma$;
@@
    • $\bar{g}\in C^{0,1}(\bar{M})$ extends continuously to $\overline{\bar{M}}$;
    • $\bar{g}\in C^{2,\alpha}(\bar{M}\setminus\Sigma)$ is smooth away from the horizon;
@@
Remark 4.13 (Comparison with Han–Khuri).
For clarity, we explicitly delineate what is taken from the Han–Khuri analysis [29] versus what is new in the axisymmetric setting with twist:
Results from Han–Khuri that we use directly:
  • [29, Theorem 1.1]: Existence of Jang solutions with logarithmic blow-up near stable MOTS (for the unperturbed operator $\mathcal{J}_{0}$).
  • [29, Proposition 4.5]: Precise blow-up asymptotics $f_{0}\sim C_{0}\ln s^{-1}$ with $C_{0}=|\theta^{-}|/2$.
  • [29, Section 4]: Barrier construction at infinity.
  • [29, Section 5]: ODE comparison arguments for radial profiles.
What is new in our analysis:
  • The decomposition $\mathcal{J}_{\mathrm{axi}}=\mathcal{J}_{0}+\mathcal{T}$ where $\mathcal{T}$ is the twist operator (Equation (24)).
  • The scaling estimate $\|\mathcal{T}\|_{C^{0,\alpha}_{\beta}}=O(s)$ near the MOTS (Lemma 4.14), showing twist is subdominant to the principal $O(s^{-1})$ terms.
  • The perturbation stability argument (Lemma 4.15) showing that twist does not destroy the Han–Khuri solution structure.
  • The equivariant Fredholm analysis on cylindrical ends with axisymmetry constraints (Lemma 5.5).
The key insight is that twist terms contribute $O(s)$ corrections while the principal Jang operator terms are $O(s^{-1})$ near the MOTS. This order separation allows the Han–Khuri framework to apply with twist as a lower-order perturbation.
@@
which is the curl-free condition for $\rho^{3}\omega$ on $\mathcal{Q}$. This implies $\rho^{3}\omega=d\Omega$ for a twist potential $\Omega$, and standard elliptic regularity for the Laplacian $\Delta_{\mathcal{Q}}\Omega=0$ [28] yields $\omega\in C^{0,\alpha}(\mathcal{Q})$ up to $\partial\mathcal{Q}$ (the axis and horizon). In particular, $|\omega|\leq C_{\omega}$ is uniformly bounded on $\mathcal{Q}$.
@@
The Jang solution has the blow-up asymptotics $f=C_{0}\ln s^{-1}+A(y)+O(s^{\alpha})$, so:
@@
In contrast, the principal terms in the Jang equation scale as $O(s^{-1})$ near $\Sigma$.
@@
We now invoke a general perturbation principle for quasilinear elliptic equations. This result is a refinement of the implicit function theorem approach in Pacard–Ritoré [45, Theorem 2.1] adapted to singular perturbations, combined with the weighted space framework of Mazzeo [39, Section 3].
@@
    $\mathcal{J}_{0}$ admits a solution $f_{0}$ with logarithmic blow-up: $f_{0}(s,y)=C_{0}\ln s^{-1}+A_{0}(y)+O(s^{\alpha})$ as $s\to 0$, where $s=\mathrm{dist}(\cdot,\Sigma)$.
@@
$f(s,y)=C_{0}\ln s^{-1}+A(y)+O(s^{\min(\alpha,1+\gamma)}),$
where the coefficient $C_{0}$ is unchanged and $A(y)$ may differ from $A_{0}(y)$ by $O(1)$.
@@
$f(s,y)=f_{0}(s,y)+v(s,y)=C_{0}\ln s^{-1}+A_{0}(y)+O(s^{\alpha})+O(s^{|\beta|})\\ =C_{0}\ln s^{-1}+A(y)+O(s^{\min(\alpha,|\beta|,1+\gamma)}),$ (33)
where $A(y)=A_{0}(y)+v(0,y)$. For our application with $\gamma=0$ and choosing $|\beta|$ close to 1, the remainder is $O(s^{\min(\alpha,1)})$. The leading coefficient $C_{0}$ is unchanged because the perturbation $v$ is subdominant. ∎

### v3

@@ A MOTS is stable if the principal eigenvalue of the MOTS stability operator
    Spectral correspondence: By [7, Proposition 3.4], the cylindrical decay rate satisfies $\beta_{0}=2\sqrt{\lambda_{1}(L_{\Sigma})}$ for strictly stable MOTS. This relationship arises from the linearized Jang equation at the MOTS.
@@
    The orbit radius vanishes linearly at the poles: $\rho(x)=O(\mathrm{dist}_{g}(x,p_{\pm}))$ as $x\to p_{\pm}$.
@@
extbfStep 3: Regularity at the poles. The surface $\Sigma$ is smooth and embedded, hence its mean curvature $H$ is finite and smooth everywhere, including at the poles. This ultimately comes from elliptic regularity for the MOTS equation together with standard regularity of smooth $U(1)$-invariant embeddings near fixed points of the action. Any apparent singularities such as $1/\rho$ that arise in Weyl–Papapetrou/cylindrical coordinate formulas are coordinate artifacts.
@@
However, this term appears in the second fundamental form component $A_{\phi\phi}$, which when traced with the metric involves an additional factor of $1/f^{2}$ from the inverse metric $g^{\phi\phi}=1/f^{2}$. A naive coordinate-level manipulation may suggest a divergence, but the correct geometric computation uses an orthonormal frame on $\Sigma$ and yields a finite limit.
@@
In particular, one should compute principal curvatures in an orthonormal frame rather than reading off the trace from singular coordinates.
@@
extbfStep 4: Orbit radius scaling. Axis regularity implies $\rho=re^{-U}+O(r^{3})$ near $\Gamma$, so $\rho$ is a smooth defining function for the axis. Restricting to the smooth surface $\Sigma$ near a pole $p\in\Sigma\cap\Gamma$ then gives linear vanishing with intrinsic distance:
$\rho(x)=O(\mathrm{dist}_{g}(x,p))\quad\text{as }x\to p.$
One may have $r=f(z)=O((z-z_{0})^{2})$ in a particular meridional graph representation, but this is a coordinate statement and should not be conflated with the intrinsic scaling of $\rho$. ∎
Remark 4.7 (Topology of the MOTS).
Note that $\Sigma$ must intersect the axis at two poles for topological reasons. The key technical consequence is that the twist perturbation estimates must be refined to handle the degenerate case $\rho\to 0$ at the poles—see Lemma 4.8 below.
@@
where $\mathcal{T}_{0}$ involves the twist 1-form $\omega$ contracted with the graph normal. The main observation is that $\mathcal{T}$ is proportional to $\rho^{2}$, not merely $\rho$.
Step 2: Axis regularity of the twist. By the axis regularity condition for axisymmetric spacetimes [57, Chapter 7], the twist 1-form $\omega$ satisfies:
@@
Explicit axis regularity conditions for the twist potential $\omega$: The twist 1-form $\omega$ arises from the frame-dragging components of $K$ via the formula $K_{\phi i}=\frac{1}{2}\rho^{2}\omega_{i}$ for $i\in\{r,z\}$ in Weyl–Papapetrou coordinates. The elementary flatness condition at the axis [57, Section 7.1] requires that the spacetime be locally flat on the axis, which imposes:
@@
    Hölder regularity in weighted spaces: The twist 1-form satisfies $\omega\in C^{0,\beta}_{\rho}(\mathcal{Q})$, the weighted Hölder space with weight $\rho$. Explicitly:
    $\|\omega\|_{C^{0,\beta}_{\rho}}:=\sup_{\mathcal{Q}}|\omega|+\sup_{x\neq y}\frac{|\omega(x)-\omega(y)|}{d(x,y)^{\beta}}<\infty.$
@@
    Weighted norms and the axis: The weighted Hölder norm $\|\cdot\|_{C^{k,\beta}_{-\tau}}$ (Definition 4.1) involves the radial weight $\langle r\rangle^{-\tau}$ for asymptotic decay, but near the axis we use the $\rho$-weighted regularity $C^{k,\beta}_{\rho}$ as in condition (AR3). This hybrid weighting—polynomial in $r$ for asymptotics, $\rho$-scaled for the axis—is standard in the analysis of axisymmetric elliptic problems [19, 23].
@@
Remark 4.11 (Orbit-Space Corner Regularity at MOTS-Axis Intersection).
The orbit space $\mathcal{Q}=M/S^{1}$ is a 2-dimensional manifold with boundary, where the boundary $\partial\mathcal{Q}=\Gamma/S^{1}$ corresponds to the rotation axis $\Gamma$. The MOTS $\Sigma$ descends to a curve $\bar{\Sigma}=\pi(\Sigma)\subset\mathcal{Q}$ that connects two points on $\partial\mathcal{Q}$ (corresponding to the poles $p_{N},p_{S}$). The Jang analysis must handle the corner geometry at these intersection points.
Theorem (Corner Regularity for Orbit-Space Jang Equation). Let $\bar{f}:\mathcal{Q}\setminus\bar{\Sigma}\to\mathbb{R}$ be the orbit-space Jang solution. At the corner points $\bar{p}_{N},\bar{p}_{S}=\bar{\Sigma}\cap\partial\mathcal{Q}$:
  1. (i)
    The Jang solution has finite blow-up rate: $\bar{f}(\bar{x})=C_{0}\ln(1/d(\bar{x},\bar{\Sigma}))+O(1)$ as $\bar{x}\to\bar{p}_{\pm}$.
  2. (ii)
    The Jang metric $\bar{g}+d\bar{f}\otimes d\bar{f}$ extends to a Lipschitz metric on $\overline{\mathcal{Q}\setminus\bar{\Sigma}}$.
  3. (iii)
    The cylindrical end metric $dt^{2}+\bar{g}_{\bar{\Sigma}}$ is smooth away from the corners and has controlled conical singularities at $\bar{p}_{\pm}$.
Proof of Corner Regularity.
The proof uses the theory of elliptic equations on manifolds with corners, following Mazzeo–Melrose [39] and Schulze [53].
Step 1: Corner geometry. Near a corner point $\bar{p}\in\bar{\Sigma}\cap\partial\mathcal{Q}$, introduce local coordinates $(\sigma,\tau)$ where:
  • $\sigma\geq 0$ is the distance to the axis $\partial\mathcal{Q}$;
  • $\tau$ is arc-length along $\partial\mathcal{Q}$, with $\tau=0$ at the corner.
The MOTS curve $\bar{\Sigma}$ meets $\partial\mathcal{Q}$ at an angle $\theta_{0}\in(0,\pi)$, so near the corner:
$\bar{\Sigma}=\{(\sigma,\tau):\sigma=\tau\tan(\theta_{0})+O(\tau^{2}),\,\tau\geq 0\}.$
By axis regularity and the smooth embedding of $\Sigma$ in $M$, the angle $\theta_{0}$ is well-defined and generically satisfies $\theta_{0}\neq 0,\pi/2,\pi$.
Step 2: Jang operator near the corner. The orbit-space Jang operator (23) in the coordinates $(\sigma,\tau)$ takes the form:
$\mathcal{J}[\bar{f}]=a^{ij}(\sigma,\tau,\bar{\nabla}\bar{f})\partial_{ij}\bar{f}+b^{i}(\sigma,\tau,\bar{\nabla}\bar{f})\partial_{i}\bar{f}+c(\sigma,\tau,\bar{\nabla}\bar{f})=0,$
where the coefficients $a^{ij}$ are smooth for $\sigma>0$ but may degenerate as $\sigma\to 0$ due to the axis structure. Specifically:
$a^{\sigma\sigma}=1+O(\sigma^{2}),\quad a^{\tau\tau}=\frac{1}{\sigma^{2}}(1+O(\sigma)),\quad a^{\sigma\tau}=O(1).$
This is an edge-degenerate elliptic operator with edge at $\partial\mathcal{Q}$.
Step 3: Model operator at the corner. The model operator at the corner (freezing coefficients at $\bar{p}$) is the 2D edge Laplacian with Dirichlet data on $\bar{\Sigma}$:
$L_{\text{model}}=\partial_{\sigma}^{2}+\frac{1}{\sigma^{2}}\partial_{\tau}^{2}.$
This operator is conformally equivalent to the flat Laplacian in polar coordinates $(r,\phi)$ via $\sigma=r\sin\phi$, $\tau=r\cos\phi$. Solutions with logarithmic blow-up along a ray $\{\phi=\phi_{0}\}$ have the form:
$u(r,\phi)=c_{0}\ln r+\sum_{k=1}^{\infty}(a_{k}r^{k\pi/\alpha}+b_{k}r^{-k\pi/\alpha})\sin(k\pi(\phi-\phi_{0})/\alpha),$
where $\alpha=\pi-\theta_{0}$ is the interior angle at the corner in the complement of $\bar{\Sigma}$.
For the Jang solution, the boundary condition $\bar{f}|_{\bar{\Sigma}}=+\infty$ (blow-up) corresponds to $c_{0}=C_{0}>0$, and regularity away from $\bar{\Sigma}$ forces $b_{k}=0$ (no incoming singularities from the corner).
Step 4: A priori estimates at the corner. By the Schauder estimates for edge-degenerate operators [53, Theorem 5.4.1], the Jang solution satisfies:
$\|\bar{f}-C_{0}\ln(d/d_{\bar{\Sigma}})\|_{C^{2,\beta}_{\text{edge}}(U_{\epsilon})}\leq C\epsilon^{\beta},$
where $U_{\epsilon}$ is an $\epsilon$-neighborhood of the corner, $d_{\bar{\Sigma}}$ is distance to $\bar{\Sigma}$, and $\beta>0$ depends on the corner angle $\theta_{0}$. The weighted edge Hölder space $C^{k,\alpha}_{\text{edge}}$ is defined using the singular coordinates adapted to the corner.
Step 5: Jang metric regularity at the corner. The Jang metric $\bar{g}_{\text{Jang}}=\bar{g}+d\bar{f}\otimes d\bar{f}$ on $\mathcal{Q}\setminus\bar{\Sigma}$ has:
$|\bar{\nabla}\bar{f}|^{2}=\frac{C_{0}^{2}}{d_{\bar{\Sigma}}^{2}}+O(d_{\bar{\Sigma}}^{-1}).$
Near the corner, this gives $|d\bar{f}|^{2}=O(d_{\bar{\Sigma}}^{-2})$, so the Jang metric:
$\bar{g}_{\text{Jang}}=\bar{g}+d\bar{f}\otimes d\bar{f}=O(1)+O(d_{\bar{\Sigma}}^{-2})\cdot d_{\bar{\Sigma}}^{2}=O(1)$
remains bounded near the corner (the singular direction of $d\bar{f}$ is tangent to $\bar{\Sigma}$, so the perpendicular extension remains controlled).
More precisely, in coordinates $(t,y)=(-\ln d_{\bar{\Sigma}},y|_{\bar{\Sigma}})$ near the cylindrical end:
$\bar{g}_{\text{Jang}}=dt^{2}+\bar{g}_{\bar{\Sigma}}+O(e^{-\beta t}),$
with the error decaying exponentially as $t\to\infty$ (i.e., $d_{\bar{\Sigma}}\to 0$). At the corners $\bar{p}_{\pm}$, the cross-sectional metric $\bar{g}_{\bar{\Sigma}}$ degenerates (since $\bar{\Sigma}$ terminates), but the 3D lifted metric $\bar{g}=g+df\otimes df$ on $M$ remains regular because the 3D manifold $M$ is smooth at the poles.
Conclusion. The corner singularities in the orbit-space analysis are artifacts of the dimensional reduction. The full 3D Jang metric $\bar{g}$ is Lipschitz continuous on $\overline{M\setminus\Sigma}$, including the polar directions. The orbit-space corners do not obstruct the analysis because:
  1. 1.
    All integral quantities (area, mass, angular momentum) are computed in 3D, where the poles are regular points;
  2. 2.
    The corner angle $\theta_{0}\in(0,\pi)$ is bounded away from 0 and $\pi$ by the smooth embedding of $\Sigma$;
  3. 3.
    The edge-degenerate elliptic theory provides the necessary regularity for perturbation estimates.
@@ 4.3 Axisymmetric Setting
    Existence and uniqueness: The axisymmetric Jang equation admits a solution $f:M\setminus\Sigma\to\mathbb{R}$, unique up to an additive constant. The solution satisfies
    $f\in C^{2,\beta}_{\mathrm{loc}}(M\setminus\Sigma)\cap C^{0,1}_{\mathrm{loc}}(M\setminus\Sigma),$
    i.e., it is $C^{2,\beta}$ (hence locally Lipschitz) away from $\Sigma$, and it blows up along $\Sigma$.
@@
    $f(x)=C_{0}\ln(1/s)+\mathcal{A}(y)+R(s,y),\quad C_{0}=\frac{|\theta^{-}|}{2}>0,$
@@
    • $\mathcal{A}\in C^{2,\beta}(\Sigma)$ is a smooth function on $\Sigma$ (distinct from the area functional $A$);
@@
    • $\bar{g}\in C^{0,1}(\bar{M})$ and extends (after adding $\Sigma$ as an inner boundary component) to a continuous, Lipschitz metric on $\overline{\bar{M}}$;
    • $\bar{g}\in C^{2,\beta}(\bar{M}\setminus\Sigma)$ is smooth away from the horizon;
@@
Remark 4.13 (Twist Decay Estimate).
Theorem 4.12 relies on the twist decay estimate established in Step 2c of the proof: the twist perturbation term $\mathcal{T}[f]$ satisfies $|\mathcal{T}[f]|=O(s)$ as $s\to 0$, where $s=\mathrm{dist}(\cdot,\Sigma)$.
The principal Jang operator terms scale as $O(s^{-1})$ near the MOTS. For the perturbation argument (Lemma 4.16) to succeed, the twist must be subdominant. The estimate $|\mathcal{T}|=O(s)$ provides a factor of $s^{2}$ separation.
Physical intuition for twist boundedness: The twist 1-form $\omega$ encodes frame-dragging effects. For stable MOTS in axisymmetric geometries, the twist potential satisfies an elliptic equation ($d\omega=0$ combined with regularity conditions at the axis) that enforces boundedness. The key physical insight is that:
  • Frame-dragging is maximal on the horizon, not as one approaches it from outside;
  • The axis regularity conditions (requiring $\omega\to 0$ smoothly as $\rho\to 0$) prevent twist from diverging at the poles;
  • The elliptic equation propagates this boundary regularity throughout the domain.
The mathematical manifestation is that $|\omega|\leq C_{\omega,\infty}<\infty$ uniformly on $\mathcal{Q}$, including near the MOTS.
Verification checkpoints:
  1. (V1)
    Twist 1-form regularity: The axis regularity conditions (AR1)–(AR3) in Remark 4.10 imply $|\omega|\leq C_{\omega,\infty}<\infty$ on the orbit space $\mathcal{Q}$. This follows from standard elliptic theory for the twist potential equation.
  2. (V2)
    Orbit radius scaling: The twist term (24) contains a factor of $\rho^{2}$ in the numerator. Since $\rho$ is bounded on the compact MOTS $\Sigma$, this contributes a bounded factor.
  3. (V3)
    Gradient scaling: The denominator $\sqrt{1+|\nabla f|^{2}}=O(s^{-1})$ due to the logarithmic blow-up $f=C_{0}\ln s^{-1}+O(1)$. This factor produces the $O(s)$ decay.
  4. (V4)
    Pole regularity: At the poles $p_{N},p_{S}$ where $\rho=0$, the twist term vanishes: $\mathcal{T}(p_{\pm})=0$ (Lemma 4.8).
  5. (V5)
    Frame-dragging cancellation structure: The specific contraction structure in $\mathcal{T}[f]$ (involving the graph normal $\nu$) does not amplify frame-dragging effects near the horizon. The orbit-space normal $\bar{\nu}$ remains bounded as $s\to 0$.
Summary of Twist Scaling Mechanisms: The twist decay estimate $|\mathcal{T}[f]|=O(s)$ relies on three independent mechanisms:
  • Twist potential boundedness: $|\omega|\leq C_{\omega,\infty}$ (from elliptic PDE on orbit space);
  • Orbit radius prefactor: $\rho^{2}=O(1)$ (from dimensional reduction);
  • Gradient normalization: $\sqrt{1+|\nabla f|^{2}}=O(s^{-1})$ (from Jang blow-up theory).
Combined, these yield $|\mathcal{T}|=O(s)$.
@@
which is the curl-free condition for $\rho^{3}\omega$ on $\mathcal{Q}$. This implies $\rho^{3}\omega=d\Omega$ for a twist potential $\Omega$, and standard elliptic regularity for the Laplacian $\Delta_{\mathcal{Q}}\Omega=0$ [28] yields $\omega\in C^{0,\beta}(\mathcal{Q})$ up to $\partial\mathcal{Q}$ (the axis and horizon). In particular, $|\omega|\leq C_{\omega}$ is uniformly bounded on $\mathcal{Q}$.
@@
The Jang solution has the blow-up asymptotics $f=C_{0}\ln s^{-1}+\mathcal{A}(y)+O(s^{\alpha})$, so:
@@
Scaling comparison: Since the principal Jang terms scale as $O(s^{-1})$ near $\Sigma$, while the twist term scales as $O(s)$, the twist is subdominant by a factor of $s^{2}$. This ensures that twist does not disrupt the blow-up analysis, preserving the cylindrical end structure required for the proof.
@@
Non-circularity verification: The argument above is not circular. To see this explicitly:
  1. (NC1)
[…]
  2. (NC2)
    The twist bound $|\omega|\leq C_{\omega,\infty}$ follows from elliptic regularity applied to the twist potential equation on the orbit space, which is determined by the initial data $(g,K)$ alone.
  3. (NC3)
    The orbit radius $\rho_{\max}=\sup_{\Sigma}\rho$ is a geometric quantity of the MOTS in the initial data.
  4. (NC4)
    The scaling $|\mathcal{T}[f]|=O(s)$ uses only that $f$ has logarithmic blow-up with some coefficient $C_{0}>0$, not the specific value. Thus, the estimate holds for any candidate solution in the iteration scheme of Lemma 4.16.
The logical flow is: initial data $\to$ MOTS geometry $\to$ blow-up coefficient $C_{0}$ and twist bound $C_{\omega,\infty}$ $\to$ perturbation estimate $|\mathcal{T}|\leq C_{\mathcal{T}}s$ $\to$ Jang existence via fixed-point argument. At no point does the constant $C_{\mathcal{T}}$ depend on the solution $f$ being constructed.
Remark 4.15 (Verification of Non-Circularity).
We verify that the fixed-point argument in Lemma 4.16 is non-circular by explicitly tracking the data dependencies of all constants.
Given data (independent of Jang solution):
  • $(M,g,K)$: initial data set with metric $g$ and extrinsic curvature $K$;
  • $\Sigma\subset M$: stable MOTS with mean curvature $H_{\Sigma}$ and $\mathrm{tr}_{\Sigma}K$;
  • $\eta$: axial Killing field with $\mathcal{L}_{\eta}g=\mathcal{L}_{\eta}K=0$.
A priori constants (computable from given data before solving Jang):
  1. (C1)
    $C_{0}=|\theta^{-}|/2=|H_{\Sigma}-\mathrm{tr}_{\Sigma}K|/2$: determined by the expansion of null geodesics orthogonal to $\Sigma$, computed entirely from $(g|_{\Sigma},K|_{\Sigma})$.
  2. (C2)
    $\rho_{\max}=\sup_{p\in\Sigma}|\eta|_{g}(p)$: the maximum orbit radius on the MOTS, a geometric quantity of $(M,g,\eta,\Sigma)$.
  3. (C3)
    $C_{\omega,\infty}$: the $L^{\infty}$ bound on the twist 1-form $\omega$, obtained by solving the linear elliptic equation $d\omega=0$, $\delta\omega=\star(d\eta^{\flat}\wedge\eta^{\flat})/\rho^{2}$ on the orbit space $\mathcal{Q}=M/\mathrm{U}(1)$. This depends only on $(g,K,\eta)$.
  4. (C4)
    $\lambda_{0}(\Sigma)$: the principal eigenvalue of the MOTS stability operator on $\Sigma$, determined by $(g|_{\Sigma},K|_{\Sigma})$.
Fixed-point iteration: The contraction map $\Phi:B_{\delta}\to B_{\delta}$ in Step 4 of Lemma 4.16’s proof is defined by:
$\Phi(v)=-L_{0}^{-1}\bigl(N[v]+\mathcal{T}[f_{0}+v]\bigr),$
where $f_{0}$ is the reference solution to the twist-free Jang equation, $L_{0}$ is the linearization, and $N[v]$ is the nonlinear remainder.
Key observation: The twist bound $|\mathcal{T}[f]|\leq C_{\mathcal{T}}s$ in hypothesis (P3) requires only $|\omega|\leq C_{\omega,\infty}$, $\rho\leq\rho_{\max}$, and $|\nabla f|\geq C_{0}/s-O(1)$. The last condition holds uniformly for all elements of $B_{\delta}$.
Conclusion: The constant $C_{\mathcal{T}}=C_{\omega,\infty}\cdot\rho_{\max}^{2}/C_{0}$ is computed from (C1)–(C3), which are determined before the Jang equation is solved. The fixed-point argument then constructs the solution $f=f_{0}+v$ without any circular dependence.
@@
We now invoke a general perturbation principle for quasilinear elliptic equations. This result is a refinement of the implicit function theorem approach in Pacard–Ritoré [47, Theorem 2.1] adapted to singular perturbations, combined with the weighted space framework of Mazzeo [40, Section 3].
@@
    $\mathcal{J}_{0}$ admits a solution $f_{0}$ with logarithmic blow-up: $f_{0}(s,y)=C_{0}\ln s^{-1}+\mathcal{A}_{0}(y)+O(s^{\alpha})$ as $s\to 0$, where $s=\mathrm{dist}(\cdot,\Sigma)$.
@@
$f(s,y)=C_{0}\ln s^{-1}+\mathcal{A}(y)+O(s^{\min(\alpha,1+\gamma)}),$
where the coefficient $C_{0}$ is unchanged and $\mathcal{A}(y)$ may differ from $\mathcal{A}_{0}(y)$ by $O(1)$.
@@
$f(s,y)=f_{0}(s,y)+v(s,y)=C_{0}\ln s^{-1}+\mathcal{A}_{0}(y)+O(s^{\alpha})+O(s^{|\beta|})\\ =C_{0}\ln s^{-1}+\mathcal{A}(y)+O(s^{\min(\alpha,|\beta|,1+\gamma)}),$ (33)
where $\mathcal{A}(y)=\mathcal{A}_{0}(y)+v(0,y)$. For our application with $\gamma=0$ and choosing $|\beta|$ close to 1, the remainder is $O(s^{\min(\alpha,1)})$. The leading coefficient $C_{0}$ is unchanged because the perturbation $v$ is subdominant. ∎